the function f is contiuous on the closed interval [0,2]

the closed interval [0,2] and has the values given in the table below. The equation f(x)=1/2 must have at least two solutions on the interval [0,2] if k= 0,1,2,3, or1/2?
Answer #1

the closed interval [0,2] and has the values given in the table below. The equation f(x)=1/2 must have at least two solutionson the interval [0,2] if k= 0,1,2,3, or 1/2?


Take the function f(x) = 1/2

Notice that it always equals 1/2

In order for it to have more than two solutions, k MUST equal 1/2. Since f(x) = 1/2 no matter what, the only "multiple" solutions arestill 1/2.

k = 1/2

answered by: Chantae